# PSDiffusion: Harmonized Multi-Layer Image Generation via Layout and Appearance Alignment

### v1

@@ 3.2. PSDiffusion
Existing multi-layer image generation models contend with inherent challenges in synthesizing coherent compositions and fostering harmonious interactions between distinct high-quality transparent foreground layers and the background image (Tudosiu et al., 2024; Zhang and Agrawala, 2024). However, it has been observed that pretrained image foundation models demonstrate proficiency in generating intricate and visually coherent compositions, effectively integrating various objects and backgrounds within a single RGB image, owing to their training on extensive datasets derived from real-world imagery. Our primary insight is to exploit the layout and interaction priors from the denoising process of the global RGB image, to inform the generation of transparent foreground layers and background layer.
@@
To fully take advantage of the layout arrangement and appearance harmonization of the global RGB images, we propose a novel framework, denoted as PSDiffusion, which concurrently generates multiple layers in a single feedforward pass, ensuring both reasonable layouts and harmonized interactions among the layers. As delineated in Figure 2, PSDiffusion comprises three principal components. First, PSDiffusion finetunes a pretrained RGB image Variational Autoencoder (VAE) into an RGBA VAE, thereby facilitating the generation of the alpha channel requisite for RGBA images, as articulated by LayerDiffuse (Zhang and Agrawala, 2024). Second, we introduce a layer cross-attention reweighting mechanism that extracts spatial layout information from the global image containing comprehensive content, generated via a global prompting strategy. This design paradigm ensures a coherent arrangement of disparate foreground objects. Finally, to promote harmonious interactions among the various foreground layers, we implement a partial joint self-attention module for inter-layer context modeling. Detailed descriptions of these two modules are provided as follows.
Figure 3. The overview of layer cross-attention reweighting module and partial joint self-attention mechanism.
@@
Layer Cross-Attention Reweighting. To ensure a reasonable layout arrangement of foreground layers, we propose a novel approach that diverges from recursive layer generation. Our key insight is to refer to the layout arrangement from the denoising process of a complete RGB image generated using the global prompt. As illustrated in Figure 3, we utilize the cross-attention block of pre-trained image foundation model to extract the layout priors. Leveraging the emergent correspondence between prompts and features—a phenomenon widely demonstrated and utilized in existing literature (Hertz et al., [n. d.]; Tang et al., 2023; Cao et al., 2023; Epstein et al., 2023; Hu et al., 2024b)—we can precisely derive and control the location of foreground instance by calculating and manipulating the cross-attention map associated with the foreground instance token. As is highlighted in the prompts in Figure 3, we select the token of the noun representing the foreground subject because its attention map most prominently displays the location of the foreground subject. Specifically, at timestep $t$ in the $n$-th cross-attention block, the global text embedding $\psi(p)$ is projected to a key matrix $\mathbf{K}^{n}=\ell_{\mathbf{K}}^{n}(\psi(p))$ and a value matrix $\mathbf{V}^{n}=\ell_{\mathbf{V}}^{n}(\psi(p))$ , and the deep spatial features of the noisy image $\phi(z_{t})$ is projected to a query matrix $\mathbf{Q}^{n}=\ell_{\mathbf{Q}}^{n}(\phi(z^{t}))$, via learned linear projections $\ell_{\mathbf{K}}^{n}$, $\ell_{\mathbf{V}}^{n}$, and $\ell_{\mathbf{Q}}^{n}$, respectively. The global attention map $\mathbf{M}^{\mathrm{g}}\in\mathbb{R}^{hw\times s}$can be calculated by
(3) $\mathbf{M}^{\mathrm{g}}=\mathrm{Softmax}\left(\dfrac{\mathbf{Q}^{n}\left(\mathbf{K}^{n}\right)^{\mathrm{T}}}{\sqrt{d}}\right).$
@@
We extend our method to obtain the layer attention map $M^{l}\in\mathbb{R}^{hw\times s}$ in the latent feature of the transparent image layer using an analogous calculation approach, where $hw$ denotes the size of the latent feature and $s$ denotes the length of the tokens. We extract the specific attention-map of the token representing the foreground subject from the global map and layer map, denoted as $\mathbf{M}_{i}^{\mathrm{g}}\in\mathbb{R}^{hw}$ and $\mathbf{M}_{i}^{\mathrm{l}}\in\mathbb{R}^{hw}$. Then implement the layer cross attention map reweighting by
(4) ${\tilde{\mathbf{M}}_{i}^{l}}$ $=\mathrm{Norm}\left(\mathbf{M}_{i}^{\mathrm{g}}\right)*\mathbf{M}_{i}^{l},$
(5) ${\hat{\mathbf{M}}_{i}^{l}}$ $={\tilde{\mathbf{M}}_{i}^{l}}/\mathrm{max}\left(\tilde{\mathbf{M}}_{i}^{l}\right)*\mathrm{max}\left(\mathbf{M}_{i}^{l}\right),$
@@
where $\mathrm{Norm}(\cdot)$ denotes Min-Max normalization on the whole feature dimension and $\max(\cdot)$ is applied on the feature dimension as well. This process provides a more precise attention map that clearly delineates the position and shape of areas associated with the foreground layer prompt. The refined attention mask is then strategically integrated into the self-attention blocks of the latent diffusion model. Through this approach, we achieve a more nuanced and controlled layout arrangement for multi-layer image generation.
@@
Partial Joint Self-attention. While our proposed layer cross-attention reweighting mechanism ensures a reasonable layout arrangement for each layer, challenges remain in harmonizing the color, style and interactions across different layers. To address the appearance inconsistency among layers, we introduce a novel partial joint self-attention mechanism for cross-layer feature context modeling. This approach facilitates more nuanced interactions among layers, ultimately achieving consistent color and lighting conditions within the composite image. As illustrated in Figure 3, we concatenate the features from global branch into the $n$-th self-attention block of the layer branch,
(6) $\mathbf{Q}=\mathbf{Q}_{L}=\mathbf{W}^{n}_{\mathbf{Q}}z_{i},\ \mathbf{K}=\mathbf{W}^{n}_{\mathbf{K}}\left[\mathbf{K}_{GL},\mathbf{K}_{L}\right],\ \mathbf{V}=\mathbf{W}^{n}_{\mathbf{V}}\left[\mathbf{V}_{GL},\mathbf{V}_{L}\right],$
@@
where $z_{i}$ denotes the latent feature for the $i$-th transparent layer, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation. $\mathbf{K}_{GL}$ and $\mathbf{V}_{GL}$ is calculated by
(7) $\mathbf{K}_{GL}=\{\mathbf{K}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\},\ \mathbf{V}_{GL}=\{\mathbf{V}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\}.$
@@
Overall, such a design enables the model to refer the appearance of the global layer, promoting a more cohesive and harmonized image generation across different foreground transparent layers and background layer.
Figure 4. Examples of our proposed inter-layer dataset.
3.3. Inter-Layer Dataset
To address the limitations of existing multi-layer datasets, we developed the dataset Inter-Layer through a human-centric curation pipeline. This dataset comprises 30,000 high-fidelity multi-layer compositions, each containing 3 - 6 layers with professional-grade alpha mattes and sophisticated inter-layer relationships. Recognizing the inadequacy of automated synthesis for achieving precise transparency channels and contextual coherence, we engaged professional designers in a multi-stage workflow. Curators systematically aggregated assets from diverse sources including web repositories, BG20K (Li et al., 2022), MAGICK (Burgert et al., 2024), and LayerDiffuse-generated transparent assets (Zhang and Agrawala, 2024), followed by rigorous manual refinement. Designers optimized each composition through: (1) Spatial layout organization with an aesthetic consideration; (2) Physics-plausible interactions among all layers; (3) Global visual harmonization across all elements. This human-in-the-loop approach ensures both technical precision in alpha channel quality and visual consistency in multi-layer storytelling, significantly advancing beyond existing datasets’ capabilities in layer complexity and visual authenticity. Some figures in our dataset are shown in Figure 4.
Figure 5. Qualitative comparison with state-of-the-art methods. Compared to existing methods, our model synthesizes multi-layer images with reasonable layout arrangement, harmonious appearance with visual effects and high alpha quality. Zoom in to check the details of foreground and composite images like shadows and edges. As exemplified by the multi-layer ensemble in the upper-right quadrant, the shadows of the elephant, man and woman are consistent with the background in their respective positions.
4. Experiments
4.1. Implementation Details
To better preserve the interaction priors of existing models, PSDiffusion is fine-tuned from Stable Diffusion XL (Podell et al., 2024) using LoRA (Hu et al., 2022). For global layer, we trained the additional $\mathbf{K}_{GL}$,$\mathbf{V}_{GL}$ linear layers in all self-attention blocks. For foreground layers and background layers, we trained $\mathbf{Q}$,$\mathbf{K}$,$\mathbf{V}$,$\mathbf{Out}$ linear layers in all attention blocks. Layer Cross-Attention Reweighting is implemented in layers at the resolution of 16, and the Partial Joint Self-attention is applied to all self-attention layers. We use the pre-trained VAE from Stable Diffusion XL for the global and background branch, and fine-tune it into an RGBA VAE to support the alpha channel of RGBA foreground image following LayerDiffuse (Zhang and Agrawala, 2024). During training, all layers of a multi-layer sample are encoded into latent space and added with the same noise at the same timestep. We use the AdamW optimizer and set the learning rate at $-41e\text{-}$ 4. Training is conducted with a total batch size of 8 on 20 NVIDIA A800 GPU hours for 10,000 iterations.
4.2. Compared with State-of-the-Art
Composite Image Layer Image

### v2

@@ 3.2 PSDiffusion
Existing multi-layer image generation models contend with inherent challenges in synthesizing coherent compositions and fostering harmonious interactions between distinct high-quality transparent foreground layers and the background image [tudosiu2024mulan, zhang2024transparent]. However, it has been observed that pretrained image foundation models demonstrate proficiency in generating intricate and visually coherent compositions, effectively integrating various objects and backgrounds within a single RGB image, owing to their training on extensive datasets derived from real-world imagery. Our primary insight is to exploit the layout and interaction priors from the denoising process of the global RGB image, to guide the generation of foreground and background layers.
@@
To fully take advantage of the layout arrangement and appearance harmonization of the global RGB models, we propose a novel framework, denoted as PSDiffusion, which concurrently generates multiple layers in a single feedforward pass, ensuring both reasonable layouts and harmonized interactions among the layers. As delineated in Figure 3, PSDiffusion comprises three principal components. First, PSDiffusion finetunes a pretrained RGB image Variational Autoencoder (VAE) into an RGBA VAE, thereby facilitating the generation of the alpha channel requisite for RGBA images, as articulated by LayerDiffuse [zhang2024transparent]. Second, we introduce a layer cross-attention reweighting mechanism that extracts spatial layout information from the global image containing comprehensive content, generated via a global prompting strategy. This design paradigm ensures a coherent arrangement of disparate foreground objects. Finally, to promote harmonious interactions among the various foreground layers, we implement a partial joint self-attention module for inter-layer context modeling.
Figure 4: Overview of our global-layer interactive mechanism, composed of layer cross-attention reweighting module and partial joint self-attention module. Layer cross-attention reweighting module extracts the cross-attention map from the global branch to reweight the cross-attention map of the layer branch, guiding the position of foreground layers. Partial joint self-attention module implements a shared attention across global branch and layer branch to facilitate context-aware feature modeling.
@@
Layer Cross-Attention Reweighting. To ensure a reasonable layout arrangement of foreground layers, we propose a novel approach that diverges from recursive layer generation. Our key insight is to refer to the layout arrangement from the denoising process of a complete RGB image generated using the global prompt. As illustrated in Figure 4, we utilize the cross-attention block of pre-trained image foundation model to extract the layout priors. Leveraging the emergent correspondence between prompts and features—a phenomenon widely demonstrated and utilized in existing literature [hertzprompt, tang2023emergent, cao_2023_masactrl, epstein2023diffusion, hu2023anomalydiffusion]—we can precisely derive and control the location of foreground instance by calculating and manipulating the cross-attention map associated with the foreground instance token. As shown in Figure 3, we select the token of the colored noun representing the foreground subject because its attention map most prominently displays the location of the foreground subject.
At timestep $t$ in the $n$-th cross-attention block, the global text embedding $\psi(p)$ is projected to a key matrix $\mathbf{K}^{n}=\ell_{\mathbf{K}}^{n}(\psi(p))$ and a value matrix $\mathbf{V}^{n}=\ell_{\mathbf{V}}^{n}(\psi(p))$ , and the deep spatial features of the noisy image $\phi(z_{t})$ is projected to a query matrix $\mathbf{Q}^{n}=\ell_{\mathbf{Q}}^{n}(\phi(z^{t}))$, via learned linear projections $\ell_{\mathbf{K}}^{n}$, $\ell_{\mathbf{V}}^{n}$, and $\ell_{\mathbf{Q}}^{n}$, respectively. The global attention map $\mathbf{M}^{\mathrm{g}}\in\mathbb{R}^{hw\times s}$can be calculated by
$\mathbf{M}^{\mathrm{g}}=\mathrm{Softmax}\left(\dfrac{\mathbf{Q}^{n}\left(\mathbf{K}^{n}\right)^{\mathrm{T}}}{\sqrt{d}}\right).$ (2)
@@
We extend our method to obtain the layer attention map $M^{l}\in\mathbb{R}^{hw\times s}$ in the latent feature of the transparent image layer using an analogous calculation approach, where $hw$ denotes the size of the latent feature and $s$ denotes the length of the tokens. We extract the specific attention-map of the token representing the foreground subject from the global map and layer map, denoted as $\mathbf{M}_{i}^{\mathrm{g}}\in\mathbb{R}^{hw}$ and $\mathbf{M}_{i}^{\mathrm{l}}\in\mathbb{R}^{hw}$. Since there are differences between the global map $\mathbf{M}_{i}^{\mathrm{g}}$ and layer map $\mathbf{M}_{i}^{\mathrm{l}}$ during the original denoising process, we utilize the global map as a guide to reweight the layer map. We implement this layer cross attention map reweighting by
$\tilde{\mathbf{M}}_{i}^{\mathrm{l}}=\mathrm{Norm}\left(\mathbf{M}_{i}^{\mathrm{g}}\right)\odot\mathbf{M}_{i}^{\mathrm{l}},$ (3)
$\hat{\mathbf{M}}_{i}^{\mathrm{l}}=\tilde{\mathbf{M}}_{i}^{\mathrm{l}}\cdot\dfrac{\max\left(\mathbf{M}_{i}^{\mathrm{l}}\right)}{\max\left(\tilde{\mathbf{M}}_{i}^{\mathrm{l}}\right)}.$ (4)
@@
where $\mathrm{Norm}(\cdot)$ denotes Min-Max normalization on the whole feature dimension, $\odot$ denotes Hadamard Product, $\max(\cdot)$ is applied on the feature dimension as well for stability,. This process provides a more precise attention map that clearly delineates the position and shape of areas associated with the foreground layer prompt. The global attention map is then strategically processed and integrated into the self-attention blocks of the latent diffusion model. Through this approach, we achieve a more nuanced and controlled layout arrangement for multi-layer image generation.
@@
Partial Joint Self-attention. While our proposed layer cross-attention reweighting mechanism ensures a reasonable layout arrangement for each layer, challenges remain in harmonizing the color, style and physics interactions across different layers. To address the appearance inconsistency among layers, we introduce a novel partial joint self-attention mechanism for cross-layer feature context modeling. This approach facilitates more nuanced interactions among layers, ultimately achieving consistent color, geometry and lighting conditions within the composite image. As illustrated in Figure 4, we concatenate the features of the global branch into the $n$-th self-attention block of the layer branch,
$\mathbf{Q}=\mathbf{Q}_{L}=\mathbf{W}^{n}_{\mathbf{Q}}z_{i}$ (5)
$\mathbf{K}=\mathbf{W}^{n}_{\mathbf{K}}\left[\mathbf{K}_{GL},\mathbf{K}_{L}\right],\ \mathbf{V}=\mathbf{W}^{n}_{\mathbf{V}}\left[\mathbf{V}_{GL},\mathbf{V}_{L}\right]$ (6)
@@
where $z_{i}$ denotes the latent feature for the $i$-th transparent layer, $\cdot$$\cdot$] denotes the concatenation operation. $\mathbf{K}_{GL}$ and $\mathbf{V}_{GL}$ are calculated by
$\mathbf{K}_{GL}=\{\mathbf{K}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\},\ \mathbf{V}_{GL}=\{\mathbf{V}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\}.$ (7)
where $\mathrm{Mask}_{i}^{\mathrm{g}}$ is derived from the global attention map $\mathbf{M}_{i}^{\mathrm{g}}$ in the Layer Cross-Attention Reweighting block and applied with filtering, blurring, and binarization, $\{\cdot\}$ denotes that we apply the attention mask in the multi-head attention. With these masks, we can restrict the $i$-th layer to query context information only from the corresponding and nearby regions in the global image, thus avoiding layer entanglement issues. Meanwhile, due to the inherent appearance harmony and element interactions of the global image, the content sharing from the associate regions of global branch to layer branch is enough for the composite harmonization.
Overall, such a design enables the model to refer to the appearance features of the global layer, promoting a more cohesive and harmonized image generation across different foreground transparent layers and background layer.
Figure 5: Examples of our proposed Inter-Layer dataset.
3.3 Inter-Layer Dataset
To address the limitations of existing multi-layer datasets, we developed the dataset Inter-Layer through a human-centric curation pipeline. This dataset comprises 30K high-fidelity multi-layer compositions, each containing 3 - 6 layers with professional-grade alpha mattes and sophisticated inter-layer relationships. Recognizing the inadequacy of automated synthesis for achieving precise transparency channels and contextual coherence, we engaged professional designers in a multi-stage workflow. Curators systematically curated assets from diverse sources including web repositories, BG20K [li2022bridging], MAGICK [Burgert_2024_CVPR], LayerDiffuse-generated transparent assets [zhang2024transparent] and Flux-generated assets [flux2024], followed by rigorous manual refinement. Designers optimized each composition through: (1) Spatial layout organization with an aesthetic consideration; (2) Physics-plausible interactions among all layers, especially geometry and orientation consistency. Designers strategically select layer combinations from our extensive RGBA asset library whose spatial and geometry relationships are pre-aligned. Minor 2D rotational and shape adjustments are then applied to optimize spatial coherence and guarantee physics-plausible contacts. (3) Global visual harmonization of all elements. Designers perform comprehensive visual adjustments to maintain consistency between foreground objects and their contextual environment. This includes calibrating saturation, hue, contrast, resolution, sharpness and brightness to match surrounding conditions. Additionally, shadows and reflections are manually crafted according to environmental factors (e.g., water surfaces, lighting direction) to achieve photorealistic integration. This human-in-the-loop approach ensures both technical precision in alpha channel quality and visual consistency in multi-layer storytelling, significantly advancing beyond existing datasets’ capabilities in layer complexity and visual authenticity. Figure 5 presents examples from our proposed dataset, highlighting its high quality, diverse multi-foreground layers, harmonized compositions, and consistent appearance.
Figure 6: Qualitative comparison with state-of-the-art methods. Compared to existing methods, our model synthesizes multi-layer images with reasonable layout arrangement, harmonious appearance with visual effects and high alpha quality. Zoom in to check the details of foreground and composite images like shadows, reflections and foreground edges. As exemplified by the multi-layer ensemble in the upper-right quadrant, the shadows of the elephant, man and woman are consistent with the background in their respective positions.
4 Experiments
4.1 Implementation Details
To better preserve the interaction priors of existing models, PSDiffusion is fine-tuned from Stable Diffusion XL [podell2023sdxl] using LoRA [hu2022lora]. For global layer, we trained the additional $\mathbf{K}_{GL}$$\mathbf{V}_{GL}$ linear layers in all self-attention blocks for joint attention. For foreground layers and background layers, we trained $\mathbf{Q}$$\mathbf{K}$$\mathbf{V}$$\mathbf{Out}$ linear layers in all attention blocks. Layer Cross-Attention Reweighting is implemented in layers at the resolution of 16, and the Partial Joint Self-attention is applied to all self-attention layers. LoRA rank is set to 256. We use the pre-trained VAE from Stable Diffusion XL for the global and background branch, and fine-tune it into an RGBA VAE to support the alpha channel of RGBA foreground image following LayerDiffuse [zhang2024transparent]. During training, all layers of a multi-layer sample are encoded into latent space and added with the same noise at the same timestep. We use the AdamW optimizer and set the learning rate at $1e\text{-}4$. Training is carried out over our Inter-Layer dataset with a total batch size of 8 on 60 NVIDIA A800 GPU hours for 10,000 iterations.
4.2 Compared with State-of-the-Art
Method Composite Image Layer Image
@@ 4.2 Compared with State-of-the-Art
LayerDiffuse [zhang2024transparent] 29.77 89.12 0.265 0.214 31.25 85.85
ART [pu2025art] 29.93 96.54 0.743 0.702 128.1
PSDiffusion(ours) 31.89 87.32 0.766 0.751 31.76 83.71
Table 1: Quantitative comparison with state-of-the-art methods. We report performance on both composite images and layer images in terms of visual quality, text alignment and composite layout harmony and interaction plausibility.
We compare our model with LayerDiffuse [zhang2024transparent] and the latest work ART [pu2025art], as they are the state-of-the-art layered image synthesis methods, and the only methods that release the source code. For evaluation, we construct 5,000 sets of multi-layer prompts, each composed of three foreground prompts, one background prompt and a global prompt describing their spatial and semantic interactions. We use vision-language model [Qwen2.5-VL] to generate the prompts to establish diverse evaluation scenarios.
We conduct quantitative evaluation using multiple metrics. For image quality assessment, we measure feature distribution discrepancies between synthesized layers and reference datasets via FID [heusel2017gans] (Fréchet Inception Distance). Specifically, the RGBA layer is evaluated against the MAGICK dataset [Burgert_2024_CVPR], while the composite images are compared with the COCO dataset [lin2014microsoft]. To assess text-image alignment for each layer, we employ the CLIP Score metric [hessel2021clipscore], which quantifies the similarity between image features and text prompts. To evaluate the harmony of the global layout and the plausibility of the layer interactions, we utilize vision-language model [Qwen2.5-VL] to conduct scoring comprehensively. The definition of layout harmony includes the rationality of the global layout and the scale of each element. The definition of interaction plausibility includes physical contacts and geometry consistency among entities, coordination of color and style, and consistency of light and shadows.
Qualitative Results. In Figure 6, we present the qualitative results on text prompt multi-layer image generation. LayerDiffuse [zhang2024transparent] tends to stack all layers at the center of the image, neglecting the spatial interactions among all layers. The latest work ART showcases reasonable layout but relies on the precise bounding boxes given by users. It also exhibits the problem of imperfect alpha quality and layer entanglement. In contrast, our model displays not only high quality of RGBA layers, but also the capability to arrange the layers to reasonable scale and position, and maintaining the visual interactions among the layers. Zoom in to see the details of foreground and composite images like edges, shadows and reflections. More qualitative results of our model are shown in Figure 2.
Quantitative Results. We report quantitative results in Table 1. We compare our PSDiffusionwith the state-of-the-art multi-layer image generation models LayerDiffuse [zhang2024transparent] and ART [pu2025art] on composite images. For LayerDiffuse, we implement the proposed sequential background-to-foreground generation to gain the multi-layer images. For ART, since it relies on the bounding-boxes as input, we first apply the LLM layout generation proposed by the authors given our global prompt, and then use the generated bounding boxes and global prompts to generate the multi-layer images. As shown in the table, our PSDiffusion outperforms LayerDiffuse and ART across four metrics. The results on layout harmony and interaction plausibility directly showcase our strengths on generating multi-layer image with enhanced layer interaction and global rationality. Furthermore, for the alpha image layer, the results in Table 1 shows that our PSDiffusion also achieves excellent alpha image quality and better text alignment.
User Study. A comprehensive perceptual evaluation was conducted with 30 participants recruited from academic institutions to holistically evaluate the visual synthesis performance of multi-layer image generation models. The study encompassed totally 171 visual samples derived from 15 distinct multi-layer text prompts, with each case containing three components: foreground layers, backgrounds layers and composite outputs generated by two baseline methods alongside our proposed PSDiffusion. Each sample underwent multi-criteria evaluation using 5-point Likert scales (1: deficient, 5: exceptional) focusing on three critical dimensions: (1) semantic-textual congruence measuring prompt relevance and contextual appropriateness, (2) visual-perceptual fidelity evaluating aesthetic coherence and artifact suppression, and (3) structural clarity assessing image definition and detail preservation. The comparative evaluation results are summarized in Table 2, demonstrating that our method consistently achieves higher user ratings across than other methods.
4.3 Ablation Studies
Layer Cross-Attention Reweighting. In this module, we utilize the cross-attention map of the global layer, and implement the reweighting operation on the cross attention map of the foreground layers to control the global layout. As shown in Figure 7, without the layer attention-map reweighting, the foreground object tends to occupy the most of the image like LayerDiffuse and even shows poor quality on the alpha channel, especially for the foreground that should have been small in the composite image.
Figure 7: Visual results of ablation study for layer cross-attention reweighting. Full PSDiffusion exhibits more plausible spatial layout including postion and scale.
Partial Joint Self-Attention. In our partial joint self-attention blocks, we implement the content sharing from global layer across foreground and background layers to encourage layer interaction plausibility like the shadows. Figure 8 displays that removing this block leads to visual effects like rigid stacks of separated foreground layers hence global disharmony.
As shown in Table 3, removing either partial joint self-attention block or layer cross-attention reweighting block results in significant decline in both layout harmony and layer interactions. This showcases their entangled importance to the coherence and interaction plausibility of the composite image.
Figure 8: Visual results of ablation study for partial joint self-attention. Full PSDiffusion demonstrates more coherent appearances and visual effects (e.g., shadows).